OpenStudy (anonymous):

One positive number is 1/5 of another number. The difference between the two numbers is 260. Find the numbers. The smaller number is_________and tell larger number is__________

OpenStudy (anonymous):

x=y/5 y-x=260 solve the system

OpenStudy (anonymous):

y-y/5=260 multiply both side by 5 5y-y=260*5 4y=1300 y=1300/4=325

OpenStudy (anonymous):

x=325/5=65

OpenStudy (anonymous):

Thank you